OpenStudy (valirium):

What are the real and complex solutions of the polynomial equation? x^4 - 41x^2 = -400

OpenStudy (tkhunny):

Have you considered the Quadratic Formula?

OpenStudy (triciaal):

let x^2 = a solve quatratic then resubstitute

OpenStudy (triciaal):

a^2 - b^2 = c^2 i = sqrt of -1

OpenStudy (solomonzelman):

Example: \(\color{black}{\displaystyle x^4+2x^2-3=0 }\) --------------------------------------- We are going to substitute \(\color{black}{\displaystyle a=x^2 }\). We get, \(\color{black}{\displaystyle x^4+2x^2-3=0 }\) \(\color{black}{\displaystyle (x^2)^2+2(x^2)-3=0 }\) \(\color{black}{\displaystyle a^2+2a-3=0 }\) \(\color{black}{\displaystyle (a-1)(a+3)=0 }\) \(\color{black}{\displaystyle a=1,~-3 }\). Then, since we had \(\color{black}{\displaystyle a=x^2 }\), we are to consider two cases: \(\color{black}{\displaystyle x^2=1 }\) \(\color{black}{\displaystyle x^2=-3 }\) \(\color{black}{\displaystyle x=\pm 1 }\) \(\color{black}{\displaystyle x=\pm \sqrt{3}i }\)

OpenStudy (solomonzelman):

So for that example you would get \(\color{black}{\displaystyle x =-1,~~1,~~\sqrt{3}i,~~-\sqrt{3}i}\) .

OpenStudy (stamp):

This is a spicy one! Solution attached, use at your own discretion. The final answer implies four roots, which is consistent with our expectations considering the degree of the polynomial was four. I have not simplified the final expression.

OpenStudy (stamp):

Simplifying the solution you'll get the nice and neat x values ±4, ±5